# Epoch-based Optimistic Concurrency Control in Geo-replicated Databases

### v1

@@ Abstract.
Geo-distribution is essential for modern online applications to ensure service reliability and high availability. However, supporting high-performance serializable transactions in geo-replicated databases remains a significant challenge. This difficulty stems from the extensive over-coordination inherent in distributed atomic commitment, concurrency control, and fault-tolerance replication protocols under high network latency.
To address these challenges, we introduce Minerva, a unified distributed concurrency control designed for highly scalable multi-leader replication. Minerva employs a novel epoch-based asynchronous replication protocol that decouples data propagation from the commitment process, enabling continuous transaction replication. Optimistic concurrency control is used to allow any replicas to execute transactions concurrently and commit without coordination. In stead of aborting transactions when conflicts are detected, Minerva uses deterministic re-execution to resolve conflicts, ensuring serializability without sacrificing performance. To further enhance concurrency, we construct a conflict graph and use a maximum weight independent set algorithm to select the optimal subset of transactions for commitment, minimizing the number of re-executed transactions. Our evaluation demonstrates that Minerva significantly outperforms state-of-the-art replicated databases, achieving over $3\times$ higher throughput in scalability experiments and $2.8\times$ higher throughput during a high network latency simulation with the TPC-C benchmark.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06copyright: none
@@ 1. Introduction
Geo-replicated distributed databases are widely used in modern large-scale online applications. In this architecture, application data is replicated to multiple data centers across a continent or even globally (sovran_transactional_2011; corbettSpannerGoogleGlobally2013; s3), providing redundancy that significantly expands fault-tolerance capability beyond the failure of a single server to failures across data centers or even the Internet infrastructure of an entire region [^1] For example, Bell, one of Canada’s Internet infrastructure providers, caused a province-wide outage in 2025 due to misconfiguration (belloutage). (fboutage2021; awsoutage2025).
@@
Prominent distributed OLTP databases, such as FoundationDB (zhou2021foundationdb), TiDB (huang2020tidb), or traditional relational databases with replication enabled, such as PostgreSQL (momjian2001postgresql), and MySQL (widenius2002mysql), replicate with a primary-backup (single-leader) architecture. In this paradigm, write operations are processed exclusively by a single replica (the primary), while secondaries serve as failover backups or read-only sources. It allows simpler concurrency control and replication protocols but presents several limitations: (1) during network partitions, availability is compromised if clients cannot reach the designated primary; (2) the single primary often becomes a performance bottleneck; and (3) primary failure requires a leader change, which is time-consuming and complexity with durability guarantees (kleppmannDesigningDataIntensiveApplications2016).
@@
To fully leverage the availability advantages of geo-replication, multi-leader (active-active, multi-master) replication is a desired approach, where all replicas are able to process both read and write requests, and propagate changes to other replicas. Doing so ensures service availability failures or network partitions, as the clients can be routed to any surviving replica. Furthermore, clients can be distributed among all replicas for load balancing and low-latency access.
@@
However, achieving high performance in multi-leader transactional geo-replicated databases remains a significant challenge, due to the extensive coordination overhead required to ensure ACID properties across replicas (bailisHAT2013; bailis_coordination_2014; zhang_tapir_2018; maiyyaCNC2019; mu_consolidating_2016; thomsonCalvinFastDistributed2012; fanOceanVistaGossipbased2019). Specifically, the system must enforce strong consistency and coordinate global concurrency control to prevent conflicting updates. These synchronization protocols often require multi-round communication, where each round trip incurs significant wide-area network (WAN) latency, ranging from 80 to 100 ms across a continent (e.g., North America) to 300 ms for intercontinental connections (e.g., US to Japan) (bailisHAT2013).
@@
Consequently, multi-leader replication is mostly seen in eventually consistent data stores, such as DynamoDB (awsdynamo2012), Cassandra (lakshmanCassandraDecentralizedStructured2010a), ScyllaDB (SycallaDB), Redis (ActiveActiveGeodistributedRedis), Riak (Riak), and Anna (wuAnnaKVSAny2021b). However, this relaxation of consistency guarantees means that full ACID transactions are no longer supported. As a result, developers must carefully reason about application correctness in the presence of stale or inconsistent reads.
@@
To address these challenges, we present Minerva, a novel distributed concurrency control protocol designed to deliver high-performance and scalable transactions that guarantee serializable isolation and crash-fault tolerance within a multi-leader architecture. Minerva is a unified protocol that combines fault-tolerant replication and concurrency control. The strong consistency and multi-leader replication also guarantee atomic commitment of transactions without requiring 2-Phase Commit (2PC). Our key design innovations are as follows.
@@
We first introduce a novel batch-based asynchronous replication protocol designed to achieve high scalability under high network latency. In our approach, replicas continuously deliver transactions in batches and establish a write-quorum. Decoupled from this data path, a coordinator periodically invokes a Raft-based (diegoRaft2014) consensus process to confirm the indices of the batches to commit. The separation of consensus agreement from bulk data replication significantly enhances system throughput.
@@
To both maximize concurrency and eliminate aborts, Minerva employs a two-phased, coordination-free transaction execution model that combines optimistic concurrency control (OCC) and deterministic execution. In the first phase, transactions are executed optimistically on local database snapshots, allowing high concurrency without locking. During replication, both transaction inputs (e.g., query parameters) and the write-sets are replicated. Upon commit, replicas independently gather the transactions in the batches, deterministically validate the transactions, then re-execute the conflicting (read-write and write-write conflicts) transactions with the same input rather than aborting them, ensuring serializability. This approach is inspired by deterministic concurrency control (thomsonddbv2010; thomsonCalvinFastDistributed2012; abadiOverviewDeterministicDatabase2018; luAriaFastPractical2020). Every commit constitutes an epoch that brings the database state a new snapshot upon each epoch transition. Thus, the only coordination is at the light-weight agreement on batch indices that is run asynchronously from the replication, transaction execution and commit process.
@@
Finally, to minimize the amount of re-executed transactions, we construct a conflict graph (bernsteinconcurrency1981) representing the concurrent conflicting transactions and then solve a maximum weight independent set (MWIS) problem to get the maximum subset of non-conflicting transactions for commitment.
@@
  • We present Minerva, a distributed OLTP database that enables scalable, fault-tolerant multi-leader replication.
@@
The remainder of this paper is organized as follows: Section 2 provides the necessary background. Section 3 details the system design and the Minerva protocol. Section 4 describes the implementation of our OLTP database. Section 5 presents the experimental evaluation, and Section 6 reviews related work.
@@ 2.1. Challenges in multi-leader Replication
In a traditional primary-backup architecture, concurrency control, atomic commitment, and replication are typically handled as orthogonal components because writes to a single data item are handled by only one replica (zhang_tapir_2018). Concurrency control is only enforced on the primary, replication only ensures consistency between the primary and its followers, and atomic commit only involves the nodes that are participating in the transaction. As a result, coordination overhead remains confined to a small subset of nodes (kleppmannDesigningDataIntensiveApplications2016).
@@
Figure 1. Single-leader vs multi-leader replication. In multi-leader, the layers for AC (Atomic Commit), $C$ (Concurrency Control), and replication among R (replicas) are no longer orthogonal to each other as in primary-backup, leading to redundant communication.
@@
Second, concurrency control must be coordinated globally. With a single primary, concurrency control is localized to the primary replica. However, with multi-leader, the system must detect concurrent reads and writes to the same data items across different replicas to prevent conflicting updates. For example, standard Two-Phase Locking (2PL) requires a transaction to acquire locks on data items across multiple replicas thus needs to coordinate (bernsteinconcurrency1981; cowlingGranolaLowOverheadDistributed2012; balakrishnanTangoDistributedData2013). Similarly, OCC executes transactions concurrently but requires a synchronized validation phase to detect conflicts and agree on the set of committed transactions (kungOptimisticMethodsConcurrency1981; ceriOCC1982; bernsteinconcurrency1981). Thus, both methods rely on consensus among replicas to guarantee serializability (thomasianDistributedOptimisticConcurrency1998; ozsuPrinciplesDistributedDatabase2020).
@@ 2.2. Toward Unified Protocols
To mitigate the overheads, prior research has proposed various optimizations designed to reduce communication overhead in distributed protocols. These strategies include leveraging inconsistent replication and delegate ordering to an asynchronous layer (e.g., TAPIR (zhang_tapir_2018)), decoupling replication from atomic commitment (e.g., Ocean Vista (fanOceanVistaGossipbased2019)), and consolidating distinct protocol phases to minimize communication rounds (e.g., Janus (mu_consolidating_2016) and the C&C framework (maiyyaCNC2019)).
More recently, deterministic concurrency control systems, such as Calvin (thomsonCalvinFastDistributed2012) and Aria (luAriaFastPractical2020), have adopted the strategy of replicating transaction inputs rather than the resulting states. In this paradigm, replicas execute transactions independently but adhere to a deterministic order. This approach eliminates the need for distributed atomic commitment and concurrency control, as the pre-established ordering guarantees serializability. By amortizing the coordination overhead across batches of transactions, these systems achieve high scalability. However, this comes at the cost of the inability to support interactive transactions and the complexity in agreeing on the commit point for each batch of transactions.
@@ 2.3. Asynchronous Consensus
Another critical challenge is to design a scalable, fault-tolerant replication mechanism. Traditional consensus protocols, such as Paxos (lamportPaxosMadeSimple2001) and Raft (diegoRaft2014) are synchronous in nature: a value is proposed, agreement is reached, and the value is persisted before the system state advances, and only then can the system advance to the next value. In high-concurrency environments, this causes significant blocking, where the agreement latency for one value delays the processing of all subsequent values (giridharanautobahn2024). Specifically, while protocols like Raft may pipeline replication messages, the requirement to commit log entries in strict sequential order means that a delay in committing one entry inevitably stalls all following entries.
Recently, some Byzantine Fault Tolerant (BFT) protocols, such as Narwhal & Tusk (danezis2022narwhal) and Autobahn (giridharanautobahn2024), address this limitation by decoupling message dissemination from the agreement process. The core idea is to propagate values continuously without blocking for immediate consensus, while a separate, asynchronous process periodically establishes the global order for batches of values.
These approaches offer high throughput consensus, but they are primarily designed for BFT systems. We adapt these BFT-oriented algorithms to be more efficient in a crash fault-tolerant scenario and integrate it into Minerva’s transaction replication mechanism.
@@
The approaches discussed above address specific challenges in building scalable, geo-replicated multi-leader databases. However, they target fundamentally different architectures and prioritize distinct trade-offs.
Inspired by these predecessors, we aim to create a unified protocol that integrates concurrency control, strongly consistent replication, and atomic commitment. Minerva adopts the ”execute-first” pattern of OCC to maximize horizontal scalability and remove coordination during the initial execution. The deterministic validation and commitment phase also allows for a coordination-free commit process. Finally, we leave the only coordination to a high-throughput asynchronous consensus protocol that decouples data propagation from agreement. These design choices collectively enable Minerva to achieve high performance and scalability in geo-distributed multi-leader environments.
3. Minerva design
@@
  • A novel concurrency control mechanism that combines optimistic local execution with deterministic re-execution, thereby maximizing concurrency while minimizing abort costs under contention while being coordination-free.
@@
  • A novel graph-based conflict detection algorithm that minimizes transaction aborts during validation by modeling the optimal commit set as the MWIS problem.
@@ 3.1. System Architecture and Assumptions
Figure 2 illustrates the system architecture. Minerva is designed with geo-distributed environments in mind where each replica is deployed in a distinct data center, forming a replication cluster that maintains a shard of the database. Clients may submit transactions to any replica, which then executes the request and propagates the changes to its peers. Within a single replica, transaction execution threads and storage nodes can be scaled horizontally with more machines. For simplicity, the remainder of this paper treats each replica as a single logical unit and assumes the replication cluster conducts full database replication  [^2] If one replication cluster only contains a partition of the databases and there are multiple clusters, 2PC is required to coordinate commit for cross-partition transactions. This can be achieved using the client as the 2PC coordinator, but it is out of the scope of this work..
@@
We define a transaction as a sequence of read, write, and computational operations. A transaction’s input consists of the client request (e.g., a SQL query string or stored procedure parameters), and its result is defined as the write-set generated upon execution.
Figure 2. Architecture of Minerva, illustrating the transaction workflow across three replicas. 1 Clients submit transactions to any replica. 2 The local replica optimistically executes transactions. 3 Inputs, read-sets, and write-sets are encapsulated in TransactionRecords. 4 These records are batched and asynchronously propagated to peer replicas. 5 A periodic commitment process triggers conflict resolution and deterministic re-execution. 6 Committed results are persisted and returned to clients.
Transactions go through four phases in Minerva:
@@
    Snapshot-based Optimistic Local Execution (§ 3.2): Clients submit requests to any replica based on locality or load. The receiving replica executes the transaction speculatively using a snapshot-based OCC protocol. Operations are performed against a consistent snapshot, derived from the last database state of the last epoch, with all write results buffered in a temporary storage.
@@
    Deterministic Validation and Re-execution (§ 3.5): At the epoch boundary, each replica independently aggregates the referenced batches and performs validation to detect read-write or write-write conflicts. To minimize aborts, a deterministic algorithm identifies the maximum subset of non-conflicting transactions using the MWIS solver. Transactions excluded from the validated set are then re-executed in a using deterministic locking (thomsonCalvinFastDistributed2012). Once the results of both the validated and re-executed transactions are persisted, the database advances to a new consistent snapshot for the epoch, and the clients are acknowledged.

### v2

@@ Abstract.
Achieving high-performance transaction processing in geo-replicated OLTP databases is challenging due to the extensive over-coordination in distributed atomic commitment, concurrency control, and fault-tolerant replication protocols. To address this issue, we introduce Minerva, a unified distributed concurrency control protocol designed for highly scalable multi-leader replication. Minerva employs a novel epoch-based asynchronous replication protocol that decouples data propagation from the commitment process, enabling continuous transaction replication. Optimistic concurrency control is used to allow replicas to execute transactions concurrently and to commit without coordination. For conflict detection during validation, we construct a conflict graph and use a maximum weight independent set search algorithm to select the optimal subset of non-conflicting transactions for commitment, minimizing the number of invalid transactions. Finally, we deterministically re-execute conflicting transactions, ensuring serializability while eliminating aborts. Our evaluation demonstrates that Minerva outperforms state-of-the-art replicated databases, achieving over $3\times$ higher throughput in scalability experiments and $2.8\times$ higher throughput in a high-latency network simulation with the TPC-C benchmark.
Distributed transactions; Multi-leader replication; Deterministic concurrency control; Geo-replication
copyright: ccjournal: PACMMODjournalyear: 2026journalvolume: 4journalnumber: 2 (SIGMOD)article: 175publicationmonth: 6doi: 10.1145/3802052ccs: Information systems Parallel and distributed DBMSsccs: Information systems Distributed database transactionsccs: Information systems Database transaction processing
@@ 1. Introduction
Geo-replicated distributed databases are widely used in modern large-scale online applications. In this architecture, application data is replicated to multiple data centers across a continent or even globally (sovran_transactional_2011; corbettSpannerGoogleGlobally2013; s3), providing redundancy that significantly expands fault-tolerance capabilities, from failures of a single server to failures across data centers or even the Internet infrastructure of an entire region [^1] For example, Bell, one of Canada’s Internet infrastructure providers, caused a province-wide outage in 2025 due to misconfiguration (belloutage). (fboutage2021; awsoutage2025).
@@
Prominent distributed OLTP databases, such as FoundationDB (zhou2021foundationdb), TiDB (huang2020tidb), or traditional relational databases with replication enabled, such as PostgreSQL (momjian2001postgresql), and MySQL (widenius2002mysql), usually replicate with a primary-backup (single-leader) architecture. In this paradigm, write operations are processed exclusively by a single replica (the primary), while secondaries serve as failover backups or read-only sources. Doing so allows simpler concurrency control and replication protocols but presents several limitations: (1) during network partitions, availability is compromised if clients cannot reach the designated primary; (2) the single primary often becomes a performance bottleneck; and (3) primary failure requires a leader change, which is time-consuming and complex with durability guarantees (kleppmannDesigningDataIntensiveApplications2016).
@@
To fully leverage the availability advantages of geo-replication, multi-leader (active-active, multi-master) replication is a desired approach, where all replicas are able to process both read and write requests, and propagate changes to other replicas. Doing so ensures service availability during failures or network partitions, as the clients can be routed to any surviving replica. Furthermore, clients can be distributed among replicas for load balancing and low-latency access.
@@
However, achieving high performance in multi-leader transactional geo-replicated databases remains a significant challenge due to the extensive coordination overhead required to ensure ACID properties across replicas (bailisHAT2013; bailis_coordination_2014; zhang_tapir_2018; maiyyaCNC2019; mu_consolidating_2016; thomsonCalvinFastDistributed2012; fanOceanVistaGossipbased2019). Specifically, the system must enforce strong consistency and coordinate global concurrency control to prevent conflicting updates. These synchronization protocols often require multi-round communication, where each round trip incurs significant wide-area network (WAN) latency, ranging from 80 to 100 ms across a continent (e.g., within North America) to over 200 ms for intercontinental connections (e.g., the US to Australia) (bailisHAT2013).
@@
Consequently, multi-leader replication is mostly seen in eventually consistent data stores, such as DynamoDB (awsdynamo2012), Cassandra (lakshmanCassandraDecentralizedStructured2010a), ScyllaDB (SycallaDB), Redis (ActiveActiveGeodistributedRedis), Riak (Riak), and Anna (wuAnnaKVSAny2021b), because weaker consistency requires less coordination (bailisHAT2013). However, the relaxation of consistency guarantees means that full ACID transactions are no longer supported. As a result, developers must carefully reason about application correctness in the presence of stale or inconsistent reads.
@@
To address these challenges, we present Minerva, a novel distributed concurrency control protocol designed to deliver high-performance and scalable transactions that guarantee serializable isolation and crash-fault tolerance within a multi-leader architecture. Minerva is a unified protocol that combines fault-tolerant replication and concurrency control. The strong consistency and multi-leader replication also guarantee atomic commitment of transactions without requiring 2-Phase Commit (2PC).
@@
The key design features of Minerva are as follows. We first introduce a novel batch-based asynchronous replication protocol designed to achieve scalability under high network latency. In our approach, replicas continuously deliver transactions in batches and establish a write quorum. Decoupled from this data path, a coordinator periodically invokes a Raft-based (diegoRaft2014) consensus process to confirm the indices of the batches to commit. The separation of consensus agreement from bulk data replication significantly enhances system throughput.
@@
To both maximize concurrency and eliminate aborts, Minerva employs a two-phased, coordination-free transaction execution model that combines optimistic concurrency control (OCC) and deterministic execution. In the first phase, transactions are executed optimistically on local database snapshots, allowing high concurrency without locking. During replication, both transaction inputs (e.g., query parameters) and the write-sets are replicated. Upon commit, replicas independently gather the transactions in the batches, deterministically validate the transactions, then re-execute the conflicting (read-write and write-write conflicts) transactions with the same input rather than aborting them, ensuring serializability. This approach is inspired by deterministic concurrency control (thomsonddbv2010; thomsonCalvinFastDistributed2012; abadiOverviewDeterministicDatabase2018; luAriaFastPractical2020). Every commit constitutes an epoch that brings the database state to a new snapshot upon each epoch transition. Thus, the only coordination is the light-weight agreement on batch indices that is run asynchronously from the replication, transaction execution, and validation.
@@
Finally, to minimize the number of re-executed transactions, we construct a conflict graph (bernsteinconcurrency1981) representing the concurrent conflicting transactions and then search for a maximum weight independent set (MWIS) in the graph to get the maximum subset of non-conflicting transactions for commitment.
@@
  • We present Minerva, a distributed concurrency control protocol that enables scalable, fault-tolerant multi-leader transactions.
@@
The remainder of this paper is organized as follows: Section 2 provides the necessary background. Section 3 details Minerva’s algorithm. Section 4 describes the implementation of our OLTP database. Section 5 presents the experimental evaluation, and Section 6 reviews related work.
@@ 2.1. Challenges in Multi-leader Replication
In a traditional primary-backup architecture, concurrency control, atomic commitment, and replication are typically handled as orthogonal components because writes to a single data item are handled by only one replica (the primary or the leader) (zhang_tapir_2018). Concurrency control is only enforced on the primary, replication only ensures consistency between the primary and its followers, and atomic commit only involves the nodes that are participating in the transaction. As a result, coordination overhead remains confined to a small subset of nodes (kleppmannDesigningDataIntensiveApplications2016).
@@
Figure 1. Primary-backup vs. multi-leader replication. In multi-leader, the layers for atomic commit (AC), concurrency control ($C$), and replication among replicas (R) are no longer orthogonal to each other as in primary-backup, leading to redundant communication.
@@
Second, concurrency control must be coordinated globally. With a single primary, concurrency control is localized at the primary replica. However, with multi-leader, the system must detect concurrent reads and writes to the same data items when transactions are executed concurrently on different replicas to prevent conflicting updates. For example, standard Two-Phase Locking (2PL) requires a transaction to acquire locks on data items across multiple replicas, and thus needs to coordinate (bernsteinconcurrency1981; cowlingGranolaLowOverheadDistributed2012; balakrishnanTangoDistributedData2013). Similarly, OCC executes transactions concurrently but requires a synchronized validation phase to detect conflicts and agree on the valid transactions (kungOptimisticMethodsConcurrency1981; ceriOCC1982; bernsteinconcurrency1981). Thus, both methods rely on consensus among replicas to guarantee serializability (thomasianDistributedOptimisticConcurrency1998; ozsuPrinciplesDistributedDatabase2020).
@@ 2.2. Toward Unified Protocols
To mitigate the overheads, prior research has proposed various optimizations designed to reduce communication overhead in distributed protocols, including the sharded primary-back up pattern. These strategies include leveraging inconsistent replication and delegating ordering to an asynchronous layer (e.g., TAPIR (zhang_tapir_2018)), decoupling replication from atomic commitment (e.g., Ocean Vista (fanOceanVistaGossipbased2019)), and consolidating distinct protocol phases to minimize communication rounds (e.g., Janus (mu_consolidating_2016) and the C&C framework (maiyyaCNC2019)).
More recently, deterministic concurrency control systems, such as Calvin (thomsonCalvinFastDistributed2012) and Aria (luAriaFastPractical2020), have adopted the strategy of replicating transaction inputs rather than the resulting states. In this paradigm, replicas execute transactions independently but adhere to a deterministic order. This approach eliminates the need for distributed atomic commitment and concurrency control, as the pre-established ordering guarantees serializability and consistency. By amortizing the coordination overhead across batches of transactions, these systems achieve high scalability. However, this comes at the cost of the inability to support interactive transactions and the complexity in agreeing on the commit point for each batch of transactions.
@@ 2.3. Asynchronous Consensus
Another critical challenge is to design a scalable, fault-tolerant replication mechanism. Traditional consensus protocols, such as Paxos (lamportPaxosMadeSimple2001) and Raft (diegoRaft2014) are synchronous in nature: a value is proposed, agreement is reached, and the value is persisted before the system state advances, and only then can the system advance to the next value. In high-concurrency environments, this causes significant blocking, where the agreement latency for one value delays the replication of all subsequent values (giridharanautobahn2024).
Recently, some Byzantine Fault Tolerant (BFT) protocols, such as Narwhal & Tusk (danezis2022narwhal) and Autobahn (giridharanautobahn2024), address this limitation by decoupling message dissemination from the agreement process. The core idea is to propagate values continuously without blocking for immediate consensus, while a separate, asynchronous process periodically establishes the global order for batches of values using causal dependencies. These approaches offer high throughput consensus, but they are primarily designed for BFT systems. We adapt these BFT-oriented algorithms to be more efficient in a crash fault-tolerant scenario and integrate them into Minerva’s transaction replication mechanism.
@@
The approaches discussed above address specific challenges in building scalable, geo-replicated multi-leader databases. However, they target fundamentally different architectures and prioritize distinct trade-offs. Inspired by these predecessors, we aim to create a unified protocol that integrates concurrency control, strongly consistent replication, and atomic commitment. Minerva adopts the ”execute-first” pattern of OCC to maximize horizontal scalability and remove coordination during the initial execution. The deterministic validation and commitment phase also allows for a coordination-free commit process. Finally, we leave the only coordination to a high-throughput asynchronous consensus protocol that decouples data propagation from agreement. These design choices collectively enable Minerva to achieve high performance and scalability in geo-distributed multi-leader environments.
3. Minerva’s Design
@@
  • A two-phased concurrency control mechanism that combines optimistic local execution with deterministic re-execution, thereby maximizing concurrency while minimizing abort costs under contention while being coordination-free.
@@
  • A graph-based conflict detection algorithm that minimizes transaction aborts during validation by modeling the optimal commit set as the MWIS problem.
@@ 3.1. System Architecture and Assumptions
Minerva is designed with geo-distributed environments in mind where each replica is deployed in a distinct data center, forming a replication cluster that maintains a copy of the database, as illustrated in Figure 2. Clients may submit transactions to any replica, which then executes the request and propagates the changes to its peers. Within a single replica, transaction execution threads and storage nodes can be scaled horizontally with more machines. For simplicity, the remainder of this paper treats each replica as a single logical unit and assumes the replication cluster conducts full database replication  [^2] If one replication cluster only contains a partition of the database and there are multiple clusters, 2PC is required to coordinate commit for cross-partition transactions. This can be achieved using the client as the 2PC coordinator, but it is out of the scope of this work..
Figure 2. Architecture of Minerva, illustrating the transaction workflow across three replicas. 1 Clients submit transactions to any replica. 2 Each replica independently and optimistically executes transactions. 3 Transaction inputs, read-sets, and write-sets are encapsulated in TransactionRecords. 4 These records are batched and asynchronously propagated to peer replicas. 5 Periodically, a consensus process is used to decide on the batches to commit, followed by independent conflict resolution and deterministic re-execution. 6 Committed results are persisted and returned to clients.
@@
We define a transaction as a sequence of read, write, and computational operations. A transaction’s input consists of the client request (e.g., SQL query strings or stored procedure parameters), and its result is defined as the write-set generated upon execution. We consider one-shot transactions but will discuss the extension to interactive transactions in § 5.9. Transactions go through four phases in Minerva:
@@
    Snapshot-based Optimistic Local Execution (§ 3.2): Clients submit requests to any replica based on locality or load. The receiving replica executes the transaction speculatively using a snapshot-based OCC protocol. Operations are performed against a consistent snapshot, derived from the database state at the end of the previous epoch, with all write results buffered in a temporary state.
@@
    Deterministic Validation and Re-execution (§ 3.5): At the epoch boundary, each replica independently aggregates the referenced batches and performs validation to detect read-write or write-write conflicts. To minimize aborts, a deterministic algorithm identifies the maximum subset of non-conflicting transactions using a MWIS solver. Transactions excluded from the validated set are then re-executed using deterministic locking (thomsonCalvinFastDistributed2012). Once the results of both the validated and re-executed transactions are persisted, the database advances to a new consistent snapshot for the epoch, and the clients are acknowledged.